Question Part
Points
Submissions Used
1 2 3
0/1 0/1 1/1
2/5 2/5 1/5
Total
1/3

An x-ray tube has an applied voltage of 185 kV.(a) What is the most energetic x-ray photon it can produce? Express your answer in electron volts and joules.
Answer in eV
Answer in J
Answer #1

The potential energy is a form of energy and the potential (and therefore voltage, when differences are taken) is defined as the potential energy (or potential energy difference) per unit charge, V=E/Q.

so E=VQ

Q of electron=1.6*10^-19 C

E=185*10^3*1.6*10^-19

E=296*10^-16 J

converting J to eV

1 joule is equal to 6.24150647996E+12 MeV.

296*10^-16 J=296*10^-16 J*6.24*10^12

=1834.56*10^-4

=0.1834MeV